# Riemannian Geometry of Optimal Rebalancing in Dynamic Weight Automated Market Makers

### 09 Mar 2026 v2

@@ Appendix F Explicit computation: SLERP vs Lambert W for $N=2$
Map to the sphere: $\boldsymbol{\mathbf{\eta}}^{\mathrm{start}}=(\sqrt{w^{\mathrm{start}}},\,\sqrt{1-w^{\mathrm{start}}})$, $\boldsymbol{\mathbf{\eta}}^{\mathrm{end}}=(\sqrt{w^{\mathrm{end}}},\,\sqrt{1-w^{\mathrm{end}}})$.
$\Omega$ $=\arccos\!\left(\sqrt{w^{\mathrm{start}}w^{\mathrm{end}}}+\sqrt{(1-w^{\mathrm{start}})(1-w^{\mathrm{end}})}\right).$ (F.17)
@@
$\boldsymbol{\mathbf{\eta}}_{\mathrm{mid}}=\frac{1}{2\cos(\Omega/2)}\left(\boldsymbol{\mathbf{\eta}}^{\mathrm{start}}+\boldsymbol{\mathbf{\eta}}^{\mathrm{end}}\right).$ (F.18)
@@
$w_{\mathrm{mid}}^{\mathrm{SLERP}}$ $\approx 0.729.$ (F.19)
Compare with $w_{\mathrm{mid}}^{\mathrm{Lambert}}\approx 0.717$ from [1]. The ${\sim}1.7\%$ difference reflects higher-order terms that matter at this (large) step size. For $f\geq 10$, per-step changes are $\leq 0.04$ and the two methods agree to $0.1\%$.
Appendix G Weight trajectories and block-to-block changes
The figures below reproduce the N=3N{=}3 setup of Figures 1–2 of [1], with the addition of SLERP, to allow direct visual comparison.
Figure G.1: Weight interpolations, $N=3$. At this scale, (b) and (c) are visually indistinguishable.
@@ Appendix G Weight trajectories and block-to-block changes
Figure G.2: Block-to-block weight changes ($\boldsymbol{\mathbf{w}}(t+1)-\boldsymbol{\mathbf{w}}(t)$). Constant metric speed produces smoothly varying (not constant) weight changes in Cartesian coordinates.
@@
Appendix H Near-boundary experiments
@@
For $N=2$ with $\boldsymbol{\mathbf{w}}:(w_{\min},\,1{-}w_{\min})\to(1{-}w_{\min},\,w_{\min})$, we compare the direct Lambert W midpoint formula of [1], $m_{i}=w_{i}^{\mathrm{end}}/W\!\bigl(w_{i}^{\mathrm{end}}\cdot e\,/\,w_{i}^{\mathrm{start}}\bigr)$ (normalised), against SLERP and the brute-force L-BFGS-B optimum.
Table H.1: Single-midpoint ($f=2$) loss ratios for symmetric $N=2$ weight changes at varying boundary distance. Lambert W, which maximises the exact retention ratio, beats SLERP near the boundary; both converge for small weight changes.
$w_{\min}$ SLERP / Optimal Lambert W / Optimal Lambert W / SLERP

### 31 Mar 2026 v3

@@ Appendix G Explicit computation: SLERP vs Lambert W for $N=2$
Map to the sphere:
$\boldsymbol{\mathbf{\eta}}^{\mathrm{start}}$ $=\bigl(\sqrt{w^{\mathrm{start}}},\;\sqrt{1-w^{\mathrm{start}}}\bigr),\quad\boldsymbol{\mathbf{\eta}}^{\mathrm{end}}=\bigl(\sqrt{w^{\mathrm{end}}},\;\sqrt{1-w^{\mathrm{end}}}\bigr).$
(G.17) $\Omega$ $=\arccos\!\left(\sqrt{w^{\mathrm{start}}w^{\mathrm{end}}}+\sqrt{(1-w^{\mathrm{start}})(1-w^{\mathrm{end}})}\right).$
@@
(G.18) $\boldsymbol{\mathbf{\eta}}_{\mathrm{mid}}=\frac{1}{2\cos(\Omega/2)}\left(\boldsymbol{\mathbf{\eta}}^{\mathrm{start}}+\boldsymbol{\mathbf{\eta}}^{\mathrm{end}}\right).$
@@
(G.19) $w_{\mathrm{mid}}^{\mathrm{SLERP}}$ $\approx 0.729.$
Compare with $w_{\mathrm{mid}}^{\mathrm{Lambert}}\approx 0.717$ from [Willetts and Harrington, 2024a]. The ${\sim}1.7\%$ difference reflects higher-order terms that matter at this (large) step size. For $f\geq 10$, per-step changes are $\leq 0.04$ and the two methods agree to $0.1\%$.
Appendix H Weight trajectories and block-to-block changes
Figure H.1. Weight differences between interpolation methods. Left: SLERP minus linear ($\sim 0.03$), reflecting the non-linearity of the geodesic. Centre: SLERP minus (AM+GM)/normalise ($\sim 0.003$), an order of magnitude smaller, with an S-curve vanishing at both endpoints and the midpoint, consistent with Theorem 6. Right: brute-force optimal minus SLERP ($\sim 10^{-7}$), confirming SLERP is near-optimal.
The figures below reproduce the N=3N{=}3 setup of Figures 1–2 of [Willetts and Harrington, 2024a], with the addition of SLERP, to allow direct visual comparison.
Figure H.2. Weight interpolations, $N=3$. At this scale, (b) and (c) are visually indistinguishable.
@@ Appendix H Weight trajectories and block-to-block changes
Figure H.3. Block-to-block weight changes ($\boldsymbol{\mathbf{w}}(t+1)-\boldsymbol{\mathbf{w}}(t)$). Constant metric speed produces smoothly varying (not constant) weight changes in Cartesian coordinates.
@@
Appendix I Near-boundary experiments
@@
For $N=2$ with $\boldsymbol{\mathbf{w}}:(w_{\min},\,1{-}w_{\min})\to(1{-}w_{\min},\,w_{\min})$, we compare the direct Lambert W midpoint formula of [Willetts and Harrington, 2024a], $m_{i}=w_{i}^{\mathrm{end}}/W\!\bigl(w_{i}^{\mathrm{end}}\cdot e\,/\,w_{i}^{\mathrm{start}}\bigr)$ (normalised), against SLERP and the brute-force L-BFGS-B optimum.
Table I.1. Single-midpoint ($f=2$) loss ratios for symmetric $N=2$ weight changes at varying boundary distance. Lambert W, which maximises the exact retention ratio, beats SLERP near the boundary; both converge for small weight changes.
$w_{\min}$ SLERP/Opt. Lambert W/Opt. Lambert W/SLERP
@@
Figure I.4. $N=3$ near-boundary configuration $\boldsymbol{\mathbf{w}}:(0.01,0.01,0.98)\to(0.49,0.49,0.02)$. Left: total arbitrage loss vs. step count $f$ for each interpolation method and the brute-force L-BFGS-B optimum. Right: relative sub-optimality (ratio of method loss to brute-force optimal loss minus one). SLERP closely tracks the brute-force optimum across all $f$; linear interpolation diverges by up to ${\sim}20\%$ near the boundary.
Figure I.5. Left: total loss at $f=1000$ as a function of the smallest weight $w_{\min}$, for $N=2$ symmetric weight changes $(w_{\min},\,1{-}w_{\min})\to(1{-}w_{\min},\,w_{\min})$. Right: loss ratio relative to SLERP. SLERP’s advantage over both linear and (AM+GM)/normalise grows as weights approach the boundary.